# MoMaGen: Generating Demonstrations under Soft and Hard Constraints for Multi-Step Bimanual Mobile Manipulation

### v1

@@ Appendix B Additional Data Generation Details
B.1 Can MoMaGen augment demonstrations with articulated objects?
We investigate whether MoMaGen can handle tasks involving articulated object manipulation. To this end, we introduce the Get Bottle task, where the robot navigates to a fridge, grasps and opens the door with one arm while maintaining stability, and then retrieves a bottle inside. This task also demonstrates MoMaGen’s ability to manage occlusion, as the bottle is initially hidden until the fridge is opened. In this case, the fridge is treated as the task-relevant object for the initial subtasks, and once the door is opened, the bottle becomes the new target. Videos of the generated demonstrations for this new task are available on our website.
@@ B.4 Visualizations of Data Diversity
Figure 10, 11, and 12 extend the visualizations in Figure 4, comparing MoMaGen to SkillMimicGen and DexMimicGen across the other three tasks. As expected, the object diversity in D1 (subfigures a) induces substantially more varied robot base and end-effector trajectories (subfigures b–c) and joint configurations (subfigure d), resulting in broader state- and action-space coverage.
@@ C.2 WB-VIMA Training Details
Figure 13: Visualization of ego-centric point cloud for Pick Cup and Tidy Table, used for WB-VIMA training. We fuse the colored point-cloud from three RGB-D cameras, crop with a robot-centric bounding box, and then downsampe to 4096 points via farthest point sampling.

### v2

@@ Appendix B Additional Data Generation Details
B.1 Can MoMaGen generalize across different families of mobile manipulation tasks?
Beyond the tasks analyzed in the main paper (rigid-body pick-and-place, wiping, and stacking), we further investigate whether MoMaGen extends to other common families of mobile manipulation behaviors. In particular, we consider three representative tasks that span different “families” of mobile manipulation: (i) articulated-object interaction, (ii) pouring and particle-transfer tasks, and (iii) whole-body control. These proof-of-concept demonstrations highlight the versatility of MoMaGen in generating stable and collision-free trajectories across diverse tasks. Videos for all tasks are provided on our project website.
Get Bottle from Fridge (Articulated Object Manipulation).
The robot navigates to a fridge, grasps and opens the door with one arm, and then retrieves a bottle inside. This task also demonstrates MoMaGen’s ability to manage occlusion, as the bottle is initially hidden until the fridge is opened. For the initial subtasks, the fridge is treated as the task-relevant object; once the door is open, the target switches to the bottle.
Pour Cat Food (Pouring and Particle Interaction).
The robot grasps a tin of cat food containing rigid particles, navigates to a bowl, and performs a controlled pouring motion until all particles are transferred. This task highlights MoMaGen’s ability to generate trajectories that require coordinated container orientation, stable end-effector motion, and interaction with particle-like objects, representative of the broader family of pouring tasks.
Push Chair (Whole-Body Control).
A chair is initially positioned outside a table. The robot navigates to the back of the chair, establishes contact, and pushes it into place underneath the table. This task requires coupling the robot’s base movement with the end-effector contact direction, demonstrating that MoMaGen can generate whole-body coordination behaviors involving sustained contact and the manipulation of large objects.
@@ B.4 Visualizations of Data Diversity
Figure 13, 14, and 15 extend the visualizations in Figure 4, comparing MoMaGen to SkillMimicGen and DexMimicGen across the other three tasks. As expected, the object diversity in D1 (subfigures a) induces substantially more varied robot base and end-effector trajectories (subfigures b–c) and joint configurations (subfigure d), resulting in broader state- and action-space coverage.
@@ B.5 Compute Resources
B.6 Failure Step Analysis
Fig. 16 illustrates the distribution of failures across the individual steps of each multi-step task. For example, in the put dishes away task, failures occur more frequently during the navigation and placement steps compared to the plate-picking step.
Figure 16: Distribution of failures across task-level steps within each multi-step task.
@@ C.2 WB-VIMA Training Details
Figure 17: Visualization of ego-centric point cloud for Pick Cup and Tidy Table, used for WB-VIMA training. We fuse the colored point-cloud from three RGB-D cameras, crop with a robot-centric bounding box, and then downsample to 4096 points via farthest point sampling.